If the equation for a market demand curve is Qd = 100 – 0,5P and the equation for a market supply curve is Qs = –20 + P, the market equilibrium quantity is:


1
Expert's answer
2021-06-07T10:35:29-0400

Market is in equilibrium when Qd=Qs

"Q_d=Q_s\\100-0.5P=-20+P\\100+20=P+0.5P\\120=1.5P\\P=\frac{120}{1,5}\\P=80"


"Q_d=100-0.5(80)=60\\Q_s=-20+80=60"


Therefore the equilibrium quantity is 60 units